# Predicting O-GlcNAcylation Sites in Mammalian Proteins with Transformers and RNNs Trained with a New Loss Function

### v2

@@ 3.1 Transformer Encoders Have Good Predictive Power on This Dataset, but are Inferior to RNNs
The primary goal of this subsection is to compare transformer models with the best RNN from Seber & Braatz (2024) using the modified dataset of Wulff-Fuentes et al. (2021) (Section 2.1) and the weighted CE loss function. While our transformer models surpass models published before Seber & Braatz (2024) and achieve an F1 score equal to 24.31% and an MCC equal to 22.49% (Table 1), transformers exhibit inferior recall at the same precision level (Fig. 1) than the RNNs from Seber & Braatz (2024), leading to inferior F1 and MCC metrics. It should be noted that the work of Pokharel et al. (2023) used an ensemble of three transformer models to predict O-GlcNAcylation and obtained considerably inferior F1 and MCC metrics than the RNN of Seber & Braatz (2024) and the transformer models in this work. It is possible that transformers are an ill-suited architecture for this task, potentially due to the amount of data available, which is why we train primarily RNNs in the following sections.
Figure 1: P-R curves for O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al. (2021). The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot. Model “YinOYang” is from Gupta & Brunak (2002); its metrics were obtained by Seber & Braatz (2024). Models “RNN-#” come from Seber & Braatz (2024). Models “Transformer-15” come from this work. The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T. The numbers on the “RNN-#” and “Transformer-15” curves represent the minimum threshold for a site to be considered positive. “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
Table 1: Performance metrics (in %) for the highest F1 score point of models tested on the modified dataset of Wulff-Fuentes et al. (2021). Models “Transformer-15” come from this work. Model “RNN-225” comes from Seber & Braatz (2024). The number after the RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T. Model “YinOYang” is from Gupta & Brunak (2002); its metrics were obtained by Seber & Braatz (2024). “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
Metric Transformer-15; 5 win (This Work) Transformer-15; 10 win (This Work) Transformer-15; 15 win (This Work) Best Threshold 0.60 0.20 0.50 Recall (%) 20.54 26.06 26.02 Precision (%) 15.63 13.77 21.28 F1 Score (%) 17.75 18.02 23.41 MCC (%) 15.50 16.08 21.36
Metric Transformer-15; 20 win (This Work) RNN-225; 20 win (Seber & Braatz (2024)) YinOYang (Gupta & Brunak (2002)) No Model Best Threshold 0.50 0.60 0.60 N/A Recall (%) 30.89 35.47 13.59 100 Precision (%) 20.04 36.90 8.08 2.44 F1 Score (%) 24.31 36.17 10.13 4.76 MCC (%) 22.49 34.57 7.57 0.00
@@
A surprising result is how the performance of the transformer models does not change significantly with the hyperparameters. Using large hidden sizes or adding another linear layer after the encoders slightly degrades the performance, while increasing the number of attention heads (to 4–5) and the number of stacked encoder layers (to 4) marginally improves the model. The most surprising result is the lack of significant improvement with increasing window size. As in Seber & Braatz (2024), our models perform better with larger window sizes for sizes up to 20 (Fig. 1 and Table 1). The performance of a transformer encoder with a window size = 5 is only marginally inferior to the RNN from Seber & Braatz (2024) using the same window size. However, whereas the RNNs’ performances increase quite significantly with increasing window size, the performance of the transformers barely improves. The only significant performance increase comes from increasing the window size from 10 to 15, but barely any performance is gained from further increasing this window size to 20 (Fig. 1).
3.2 The weighted focal differentiable MCC loss function is a superior alternative to the weighted CE
We hypothesized that changing the loss function can lead to improvements in the models’ metrics, as the weighted CE is optimizing a function that is only correlated with F1 scores and MCC. Four other loss functions (as described in Section 2.4) are also used to cross-validate transformer models through the same procedure used with the weighted CE function (Sections 2.2 and 3.1).
@@ 3.2 The weighted focal differentiable MCC loss function is a superior alternative to the weighted CE
The Focal, Lovász, and differentiable F1 losses fail to achieve meaningful results. In the best-case scenarios, models trained with these loss functions can achieve F1 scores equal to or slightly above 4.74%, equivalent to treating all (or almost all) sequences as positive. The differentiable F1, in particular, frequently gets stuck at this local F1 maximum. Naturally, these models have an MCC $\approx 0$, as they are no better than random classifiers.
Conversely, the weighted focal differentiable MCC loss is able to attain very good performance on this problem (Fig. 2 and Table 2). For both architectures, its performance is superior to that of the weighted cross-entropy loss. As in Section 3.1, RNNs are considerably better than transformer encoders in predicting O-GlcNAcylation. The RNN model trained with the weighted focal differentiable MCC loss achieves an F1 score = 37.03% and an MCC = 36.58%. The relative and absolute increases in MCC are higher than those for the F1 score, highlighting the effectiveness of the weighted focal differentiable MCC loss, as it seeks to directly optimize the MCC. The weighted focal differentiable MCC loss can also be used for fine tuning; an RNN model first trained using weighted cross-entropy, then fine-tuned using the weighted focal differentiable MCC is generated (“RNN-225; CE $\rightarrow$ diff MCC”). This model initially followed the optimal hyperparameters (including model size) found through cross-validation using the weighted cross-entropy loss, but the learning rate and loss hyperparameters are changed to the optimal values found through cross-validation with the weighted focal differentiable MCC loss function when that loss function is used for fine-tuning. This fine-tuning also leads to a model that is better than the previous state-of-the-art from Seber & Braatz (2024) (Fig. 2 and Table 2), reaching an F1 score = 36.52% and an MCC = 36.01%, values slightly lower than that of the model trained only with the weighted focal differentiable MCC loss.
There is a significant difference between the optimal hyperparameters selected through the weighted cross-entropy and weighted focal differentiable MCC loss functions. Compared to the optimal values found through the weighted cross-entropy loss, the weighted focal differentiable MCC loss has an optimum at higher model hidden sizes (15 vs. 240 for transformers; 225 vs. 1425 for RNNs), higher batch sizes (256 vs. 512 for transformers; 32 vs. 128 for RNNs), lower learning rates (1×$\times 10^{-3}$ vs. 1×$\times 10^{-4}$), and lower class weights (15 vs. 3 for transformers; 30 vs. 3 for RNNs).
Figure 2: P-R curves for O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al. (2021). The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot. Model “RNN-225; 20 win; CE” comes from Seber & Braatz (2024). All others come from this work. The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T. The numbers on the curves represent the minimum threshold for a site to be considered positive. “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
@@
Table 2: Recall, precision, F1 score, and MCC metrics (in %) for the highest F1 score point of models tested on the modified dataset of Wulff-Fuentes et al. (2021). Model “RNN-225; CE” comes from Seber & Braatz (2024). All others come from this work. The number after each RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; the number after the semicolon for both types of models represents the window size on each side of the central S/T. “CE” represents the weighted cross-entropy loss, and “diff MCC” represents the weighted focal differentiable MCC loss. “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
@@
Metric Transformer-15; CE (This Work) Transformer-240; diff MCC (This Work) RNN-225; CE (Seber & Braatz (2024)) Best Threshold 0.50 0.40 0.60 Recall (%) 30.89 25.70 35.47 Precision (%) 20.04 25.48 36.90 F1 Score (%) 24.31 25.59 36.17 MCC (%) 22.49 23.67 34.57
@@
Metric RNN-1425; diff MCC (This Work) RNN-225; CE $\rightarrow$ diff MCC (This Work) No Model Best Threshold 0.70 0.40 N/A Recall (%) 30.67 30.35 100 Precision (%) 46.70 45.84 2.44 F1 Score (%) 37.03 36.52 4.76 MCC (%) 36.58 36.01 0.00
3.3 Stacked RNNs provide better performance, and the weighted focal differentiable MCC loss is still superior
Seber & Braatz (2024) trained RNNs with only a single cell, a restriction that was followed by this work in Section 3.2. However, RNNs with multiple stacked cells may perform better than single-cell RNNs. To verify this hypothesis, RNN models with 2, 3, and 4 LSTM layers are tested (as per Section 2.3). Similarly to Section 3.2, three models are created: one trained only with the weighted CE loss, one trained only with the weighted focal differentiable MCC loss, and one first trained with the weighted CE loss, then fine-tuned with the weighted focal differentiable MCC loss. These models all surpass their single-RNN layer equivalents from Section 3.2. The model trained with the weighted CE loss (“RNN-[450,75]; CE”) displays the greatest relative improvement over its single-cell version, reaching an F1 score = 38.73 % and an MCC = 37.31%. Like with the single-cell RNNs, fine-tuning with the weighted focal differentiable MCC loss after training with the weighted CE leads to improvements in the model performance; this fine-tuned model (“RNN-[450,75]; CE $\rightarrow$ diff MCC”) reaches an F1 score = 38.83% and an MCC = 37.33%. While the performance improvement of fine-tuning is smaller in this scenario, these improved results again confirm the efficacy of fine-tuning with the weighted focal differentiable MCC loss. Finally, a model trained directly with the weighted focal differentiable MCC loss (“RNN-[600,75]; diff MCC”) achieves the best results, reaching an F1 score = 38.88% and an MCC = 38.20%. The increase in MCC is considerably higher, further corroborating the benefits of this new loss function discussed in Section 3.2. Overall, the “RNN-[600,75]; diff MCC” model displays a 7.5% increase in F1 score and a 10.5% increase in MCC over the “RNN-255; CE” model of Seber & Braatz (2024).
Figure 3: P-R curves for two-layer O-GlcNAcylation site prediction models tested on the modified dataset of Wulff-Fuentes et al. (2021). The black dotted lines are isolines of the F1 score, as labeled on the top and right sides of the plot. The number after each RNN model is the LSTM module size used, and the number after the semicolon represents the window size on each side of the central S/T. The numbers on the curves represent the minimum threshold for a site to be considered positive. “No Model” sets all sites to potentially positive – that is, sets the threshold to 0.
Table 3: Recall, precision, F1 score, and MCC metrics (in %) for the highest F1 score point of two-layer models tested on the modified dataset of Wulff-Fuentes et al. (2021). The number after each RNN model is the LSTM module size used, and the number after the semicolon represents the window size on each side of the central S/T. “CE” represents the weighted cross-entropy loss, and “diff MCC” represents the weighted focal differentiable MCC loss. The “No Model” (threshold = 0) baseline metrics are as in Tables 1 and 2.
The models with the highest average cross-validation F1 score all had only two layers and second LSTM layers with 75 neurons. As in Section 3.2, the optimal hyperparameters for models trained with the weighted focal differentiable MCC loss included higher model hidden sizes (600 vs. 450 for CE), lower learning rates (5×$\times 10^{-4}$ vs. 1×$\times 10^{-3}$), and lower class weights (2 vs. 20). Surprisingly, both models had the same optimal batch size (128), and the hidden size and class weights for the model trained with the weighted CE loss moved towards those from the model trained with the weighted focal differentiable MCC loss.
Finally, we verified that our splits for the cross-validation and test sets do not bias the models by performing five-fold nested cross-validation, following the procedure in Section 2.1, on the “RNN-[600,75]; diff MCC” model. The models generated by each fold had very similar performance to each other (Fig. A1 and Table A1), highlighting the robustness of that architecture and loss function to changes in the training/test data.
4 Discussion
This work generates transformer encoder (Section 3.1) and RNN (Sections 3.2 and 3.3) models to predict O-GlcNAcylation sites based on protein sequences using a modified version of an extensive public dataset Wulff-Fuentes et al. (2021). These models are also compared to the state-of-the-art RNNs trained by Seber & Braatz (2024).
Although transformer models attain high performances on this dataset, they are inferior to the RNN models trained by Seber & Braatz (2024), reaching an F1 score of 24.31% and an MCC of 22.49% (Section 3.1, Fig. 1, and Table 1). A significant difference between transformers and RNNs is that the former do not scale well with increasing window sizes, while the latter display significant improvements with larger window sizes. With window size = 5, both transformers and RNNs possess similar performances, but a large performance gap arises with increasing window sizes. Multiple reasons could be causing this inferior transformer performance, including a lack of training data. Although the dataset from Wulff-Fuentes et al. (2021) is one of the largest (if not the largest) O-GlcNAcylation datasets published, it is possible that there simply are not enough data for the transformer architecture to perform well in this task.
Because the weighted cross-entropy loss function, which is used to train the models in Seber & Braatz (2024) and Section 3.1 of this work, only indirectly optimizes the classification performance metrics, we hypothesized that a function that could directly improve the F1 score and MCC can lead to better-performing models.
We develop the weighted focal differentiable MCC loss to address this question (Section 2.4; Algorithm 1). This function allows the direct optimization of the MCC, yielding better classification models. We cross-validate and test transformer and RNN models using the same dataset and notice that the performance of the weighted focal differentiable MCC loss is similar to but higher than that of the weighted cross-entropy loss (Section 3.2, Fig. 2, and Table 1). This loss function improves both the transformer and RNN architectures, and leads to the creation of an improved model with an F1 score = 37.03% and an MCC = 36.58%. Furthermore, a model first trained using the weighted cross-entropy loss, then fine-tuned using the weighted focal differentiable MCC loss achieves improved performance in this prediction task relative to a model trained using only the weighted CE loss, reaching an F1 score = 36.52% and an MCC = 36.01%. While these metrics are slightly smaller than the values obtained by a model directly trained with the weighted focal differentiable MCC loss, these results show how this new loss function can also be used to optimize and improve trained models at a low cost. The optimal hyperparameters for models trained with the weighted focal differentiable MCC loss are different from those of models trained with the weighted cross-entropy loss.
As Seber & Braatz (2024) used only single-cell models in their work, the RNN models from Section 3.2 follow this restriction. Models with multiple stacked cells can have better performance, so we train RNN models with multiple LSTM cells in Section 3.3. These models display superior cross-validation F1 scores over models with one cell, and the best models have only 75 neurons in their 2nd cells. No matter the loss function used, these deeper models display better performance in the test set than the best RNN of Seber & Braatz (2024) (Fig. 3 and Table 3). As in Section 3.2, fine-tuning using the weighted focal differentiable MCC loss after training with the weighted CE leads to improvements in performance over using only the weighted CE, corroborating one of the benefits of the weighted focal differentiable MCC loss. The model trained solely with the weighted focal differentiable MCC loss, “RNN-[600,75]; diff MCC”, achieves state-of-the-art performance in O-GlcNAcylation prediction, reaching an F1 score = 38.88% and an MCC = 38.20%. Overall, this is a 7.5% increase in F1 score and a 10.5% increase in MCC over the “RNN-255; CE” model of Seber & Braatz (2024).
There are two main limitations to this work. On the O-GlcNAcylation prediction side, although the models developed in this work reach state-of-the-art performance in multiple independent datasets, that performance is still far away from (near-)perfect, possibly limiting the widespread usability of the models. It should be noted that the experimental detection of O-GlcNAcylation sites is a laborious and expensive process, so even an imperfect model may be useful. Having a usable model is particularly important given the relevance of O-GlcNAcylation to many diseases and the potential of O-GlcNAcylation as a target for medicines (Section 1). On the weighted focal differential MCC loss side, although we earnestly believe that it can improve the performance of models in almost any classification task relative to the (weighted) CE, this work tests its performance on only one task. Future directions primarily include testing the weighted focal differential MCC loss on multiple classification tasks to show its wide usability, including in fine-tuning models that were previously trained with the CE loss. Another direction is analyzing the effects of dynamically changing this loss function’s hyperparameters over the epochs to determine whether it can be further improved.
The code used in this work is publicly available, allowing the reproduction, improvement, and reuse of this work. It is simple to install and run the best RNN model to predict O-GlcNAcylation sites based on the local protein sequence. Instructions are provided in Sections A1.1 and A1.2 of the supplemental data or the README in our GitHub <redacted for review> repository.

### v3

@@ Methods
The primary goal of this subsection is to compare transformer models with the best RNN from Ref. [6], labeled in this work as “RNN-225” or “RNN-225; CE”, using the modified dataset of Ref. [9] (Section Dataset). While this work’s transformer models surpass models published before Ref. [6] and achieve an F1 score equal to 24.31% and an MCC equal to 22.49% (Table S1), transformers exhibit inferior recall at the same precision level (Fig. S1) than the RNNs from Ref. [6], leading to inferior F1 and MCC metrics. It should be noted that the work of Ref. [11] used an ensemble of three transformer models to predict O-GlcNAcylation and obtained considerably inferior F1 and MCC metrics than the RNN of Ref. [6] and the transformer models in this work.
@@
A surprising result is how the performance of transformer encoder models does not change significantly with the hyperparameters used. Using large hidden sizes or adding another linear layer after the encoders slightly degrades the performance, while increasing the number of attention heads (to 4–5) and the number of stacked encoder layers (to 4) marginally improves the model. The most surprising result is the lack of significant improvement with increasing window size. As in Ref. [6], our models perform better with larger window sizes for sizes up to 20 (Fig. S1 and Table S1). The performance of a transformer encoder with a window size = 5 is only marginally inferior to the RNN from Ref. [6] using the same window size. However, whereas the RNNs’ performances increase quite significantly with increasing window size, the performance of the transformers barely improves. The only significant performance increase comes from increasing the window size from 10 to 15, but little performance is gained from further increasing this window size to 20 (Fig. S1).
@@
We hypothesized that changing the loss function can lead to improvements in the models’ metrics, as the weighted CE is optimizing a function that is only correlated with F1 scores and MCC. Four other loss functions (as described in Section Loss Functions) are also used to cross-validate transformer models through the same procedure used with the weighted CE function. The Focal, Lovász, and differentiable F1 losses fail to achieve meaningful results. In the best-case scenarios, models trained with these loss functions can achieve F1 scores equal to or slightly above 4.74%, equivalent to treating all (or almost all) sequences as positive. The differentiable F1, in particular, frequently gets stuck at this local F1 maximum. These models have an MCC $\approx 0$, as they are no better than random classifiers. However, a transformer model trained with the weighted focal differentiable MCC was able to surpass the performance of the transformers trained with the weighted CE loss (Table S1 and Fig. S1), improving the F1 score to 25.59% and the MCC to 23.67%. Nevertheless, these metrics remain inferior to the previous state-of-the-art RNN model.
@@
Table S1: Performance metrics (in %) for the highest F1 score point of models tested on the modified dataset of Ref. [9]. Models “Transformer-#” come from this work. The “RNN-225” model [6] is the previous state-of-the-art model. The number after the RNN model is the LSTM module size used; the number after each transformer model is the hidden size used; WS is the window size on each side of the central S/T. Metrics for the model “YinOYang” [7] were obtained by Ref. [6]. Baseline (threshold = 0) metrics are precision = 2.44%, F1 score = 4.76%, and MCC = 0%.
@@
Metric Transformer-15; CE (This Work) WS = 5 WS = 10 WS = 15 WS = 20 Best Threshold 0.60 0.20 0.50 0.50 Recall (%) 20.54 26.06 26.02 30.89 Precision (%) 15.63 13.77 21.28 20.04 F1 Score (%) 17.75 18.02 23.41 24.31 MCC (%) 15.50 16.08 21.36 22.49
@@
Metric Transformer-240; diff MCC (This Work) RNN-225; CE (Ref. [6]) YinOYang (Ref. [7]) WS = 20 WS = 20 Best Threshold 0.40 0.60 0.60 Recall (%) 25.70 35.47 13.59 Precision (%) 25.48 36.90 8.08 F1 Score (%) 25.59 36.17 10.13 MCC (%) 23.67 34.57 7.57
Fig S1: P-R curves for O-GlcNAcylation site prediction models tested on the modified dataset of Ref. [9]. Labels, curve descriptions, and baseline metrics are as in the captions of Table S1 and Fig. 1.
S4.2 Nested validation of the two-layer RNN models
Fig S2: Precision-recall curves for the “RNN-[600,75]; diff MCC” model after nested validation on the modified dataset of Ref. [9]. The black dotted lines are F1 score isolines, as labeled on the top and right sides of the figure.
Table S2: Recall, precision, F1 score, and MCC metrics (in %) for the point with highest F1 score for the “RNN-[600,75]; diff MCC” model after nested validation on the modified dataset of [9]. Baseline metrics are as in the caption of Table S1.
[…]
Table S3: Recall, precision, F1 score, and MCC metrics (in %) for the point with highest F1 score for the “RNN-[450,75]; CE” model after nested validation on the modified dataset of [9]. Baseline metrics are as in the caption of Table S1.
[…]
S5 Computational Resources Used
All experiments were done in a personal computer equipped with a 13th Gen Intel® Core™ i5-13600K CPU, 64 GB of DDR4 RAM, and an NVIDIA GeForce RTX 4090 GPU.